Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on November 22nd

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • MOM report
    • Started 'running' on Friday evening with a 24-hour shift.
    • ISIS very unstable. Switched on Sunday to TS2 timing.
    • Conventional water issues, now resolved.
    • Superconducting magnets regularly monitored. Very stable.
    • Run plan:
      • Have data for 2 of the settings.
      • Prepared a third beam setting for recording over night
      • More beam tuning tomorrow.
    • Loss of shift cover for SS monitoring at 4pm tomorrow.
    • Ramp magnets down at 2pm (Agreed).
  • data summary as of 3.30 pm yesterday - Rogers gives apologies
  • SS Magnet status
    • Magnet alarms need adjusting, to reflect running performance.
    • Feedback from joseph to pierrick regarding operation by non-ss personel.
  • Hydrogen status
    • Cooling, Absorber at 46K. 2nd stage ~26K.
    • Monitor overnight and determine coldest temperature. Then will try using heater.
  • Conventional magnet water flow issue
    • Removed PLC from both contentional water pumps.
    • Pump set to deliver constrant pressure from configuration app.
  • Wednesday maintenance plans
    • No plans.
  • Run plans for the next week (discuss)
    • Wednesday 23rd : ramp down the SS magnets / maintenance
    • Thursday : Shift training + 1 shift - straight track running
    • Friday - Sunday : 1 shift per day - straight track running
    • Monday 28th : ramp up SS magnets
  • Missing shifts
    • Still have some missing shift slots for minimal running
      • Friday 25th : BLOC
      • Dec 3 : Overnight monitoring shift (00:00 - 08:00)
  • AOB

Updated by Overton, Edward almost 7 years ago ยท 11 revisions